Route Canal Shipping Company has the following schedule for aging of accounts receivable:

Age of Receivable

( 1) ( 2 ) ( 3 ) ( 4 )

Month of Sales Age of Account Amounts % of amount due

April 0-30 $195,360

March 31-60 $97,680

February 61-90 $122,100

January 91-120 $73,260

Total receivables $488,400 100%

a. Calculate the percentage of amount due for each month

b. If the firm had $1,584,000 in credit sales over the four-month period, compute the average collection period. Avg. daily sales should be based on a 120-day period.

c. If the firm likes to see its bills collected in 42 days, should it be satisfied with the average collection period?

d. Disregarding your answer to part c and considering the aging schedule for accounts receivable, should the company be satisfied?

Please show your answer in a way that I can learn to solve it myself, thanks!
Answer #1

a) april=195360=488400=40%

march=97680/488400=20%

feb=122100/488400=25%

jan=73260/488400=15%


b)488400*120/1584000=37 days


c)no, its doesnt satisfies.it should be 37 days


d)company shoud not be satisfied
Answer #2

Hi,


Please find the answer as follows:


Part A:


Percentage of amount due for each mont:

April = Monthly Amount Receivables/Total Value of Receivables = 195360/488400*100 = 40%

March = Monthly Amount Receivables/Total Value of Receivables = 97680/488400*100 = 20%

February = Monthly Amount Receivables/Total Value of Receivables = 122100/488400*100 = 25%

January = Monthly Amount Receivables/Total Value of Receivables = 73260/488400*100 = 15%


Part B:

Average Collection Period = 120/Accounts Receivables Turnover Ratio

Accounts Receivables Turnover Ratio = Credit Sales/Accounts Receivables = 1584000/488400 = 3.24


Average Collection Period = 120/3.24 = 37 days.


Part C:

Since the average collection period should ideally be 37 days, the company should not be satisfied with average collection period of 42 days.


Part D:

No, the company should not be satisfied.


Thanks.